# Aligning Large Language Models with Self-generated Preference Data

### v1

@@ Abstract
Aligning large language models (LLMs) with human preferences becomes a key component to obtaining state-of-the-art performance, but it yields a huge cost to construct a large human-annotated preference dataset. To tackle this problem, we propose a new framework that boosts the alignment of LLMs through Self-generated Preference data (Selfee) using only a very small amount of human-annotated preference data. Our key idea is leveraging the human prior knowledge within the small (seed) data and progressively improving the alignment of LLM, by iteratively generating the responses and learning from them with the self-annotated preference data. To be specific, we propose to derive the preference label from the logits of LLM to explicitly extract the model’s inherent preference. Compared to the previous approaches using external reward models or implicit in-context learning, we observe that the proposed approach is significantly more effective. In addition, we introduce a noise-aware preference learning algorithm to mitigate the risk of low quality within generated preference data. Our experimental results demonstrate that the proposed framework significantly boosts the alignment of LLMs. For example, we achieve superior alignment performance on AlpacaEval 2.0 with only 3.3% of the ground-truth preference labels in the Ultrafeedback data compared to the cases using the entire data or state-of-the-art baselines.
@@ 1 Introduction
Recently, large language models (LLMs) have made huge progress in various NLP tasks, leading to real-world applications that are used by millions of users, such as coding assistants and chatbots [1, 27, 38]. Aligning LLMs with human feedback, particularly through learning from human preferences, is widely considered a crucial technique for their success [5, 18, 51]. To enhance this alignment, various preference learning algorithms have been extensively explored [29, 32]. Despite these advancements, one of the remaining challenges is the reliance on large-scale human-annotated preference data. As the quality and quantity of preference data are critical for the successful alignment of LLMs [2, 7], the huge cost to acquire such data inevitably presents significant obstacles.
@@
To mitigate this challenge, engaging LLMs in constructing preference data and improving their alignment using this data has recently gained attention. For example, a representative way on this line is generating multiple responses for the input prompts, and then approximating human preference between them through LLM’s predictions, often referred to as LLM-as-judge [3, 48]. However, these approaches are only effective when the given LLM is sufficiently large and well-aligned to mimic human preference via in-context learning. On the other hand, using an external reward model is considerable to substitute human preference annotation efficiently [16, 35], but it is built on the availability of large human preference data and also could be ineffective if the distribution mismatch exists. Lastly, these approaches have a risk of potential labeling noise from LLMs, but this aspect has not been explored yet. Therefore, in this work, we aim to develop a method to effectively improve the alignment of LLM by overcoming these limitations but only relying on the small human annotation.
@@
Figure 1: Illustration of the proposed Selfee framework. Selfee progressively improves the alignment of LLMs by iterating (1) the generation of new preference data and (2) the preference learning on the constructed data with self-refinement. Technical details are presented in Section 4.
@@
Contribution. We introduce a simple yet effective framework, coined Selfee, to improve the alignment of LLMs via Self-generated Preference data, with only a small amount of human-labeled preference data. Our key idea is to progressively expand the knowledge of human preference within the small (seed) data, by iteratively generating the responses and learning from them through the self-annotated preference labels. Specifically, our technical contributions are three-fold as described in what follows. First, we derive the preference labels using the logits of LLM to explicitly extract the model’s inherent preference. This approach is more effective compared to previous methods that rely on external reward models or implicit in-context learning. Second, we introduce a confidence-based refinement of preference labels to reduce the risk of noise in preference learning with generated data. Third, to further enhance the effectiveness of this refinement, we propose using a linearly extrapolated prediction between the current and reference models; this approach approximates the predictions of a more strongly aligned model, leading to better noise identification.
@@
Figure 2: Summary of main result. Evaluation results on AlpacaEval2.0 benchmark [20]. Our framework significantly improves the alignment of LLMs, without additional human preference data. See detailed results in Section 5.
@@
We demonstrate the effectiveness of the proposed Selfee by aligning recent LLMs with small human-annotated preference data and evaluating their alignment on the commonly used benchmarks. For example, using only 3.3% of ground-truth preference in Ultrafeedback data [7] with the mistral-7b-0.1v SFT model [15], our framework achieves over 16.4% increase in AlpacaEval2.0 [20] win rate compared to the initial SFT model (see Figure 2). Additionally, the AlpacaEval2.0 length-controlled win rate is improved from 7.58% to 15.39%, and MT-bench score [50] increased from 6.38 to 6.94. Compared to preference judgment methods like LLM-as-judge [50], and even strong reward models such as PairRM [16], which have recently shown state-of-art performance in AlpacaEval2.0 benchmark, our approach consistently outperforms them across all metrics. More interestingly, the proposed Selfee successfully improves the alignment of various LLMs, even without the initial human preference data. These results demonstrate that our framework is highly competitive and practical for real-world applications.
@@ 2 Related Work
Alignment of LLMs with human preference. Learning from human preferences now serves as a core component for the state-of-the-art LLMs [1, 28, 38, 39] for aligning their responses with users’ intent and values [29, 51]. Arguably, one of the most popular frameworks is reinforcement learning with human preference (RLHF) [5, 18], which first trains the reward model, and then fine-tunes LLM to maximize that reward with KL divergence regularization to prevent the reward over-optimization of LLM. On the other hand, various training algorithms have recently been proposed to fine-tune LLMs with human preference more efficiently [11, 14, 23, 32, 47, 49]. For example, Rafailov et al. [32] proposes Direct Preference Optimization (DPO) which allows one to fine-tune LLMs without a separate reward modeling stage, by deriving the training objective mathematically equivalent to RLHF. Ethayarajh et al. [11] further removes the reliance on pair-wise preference labels by formulating the objective based on a human utility model. However, these methods assume that large human-annotated preference data is available, which requires a huge data acquisition cost.
Engagement of LLMs for constructing preference data. For an efficient and scalable alignment procedure, engaging LLMs for preference dataset construction has recently received attention. One common approach involves generating multiple responses to input prompts from LLM, and using an LLM’s predictions to approximate human preferences between them, a technique often referred to as LLM-as-judge [2, 48]. However, this method is effective only when the LLM is sufficiently large and well-aligned to mimic human preferences through in-context learning. Alternatively, employing an external reward model can efficiently replace human preference judgment [16, 35], but this approach relies on the availability of extensive human preference data to pre-train reward model and may be ineffective if there is a distribution mismatch. Some concurrent works [34, 35, 45, 46] have proposed the alignment procedure with iterative data expansion and preference learning. However, they use the external reward model or stronger LLM for the preference judgment. In contrast, we only utilize the intrinsic knowledge of training LLM with new data expansion and preference learning methods.
@@ 3 Preliminaries
Let denote LLM as $\pi_{\theta}$, which generates an output sequence (e.g., response) $y$ for a given input sequence (e.g., prompt) $x$, i.e., $y\sim\pi_{\theta}(\cdot|x)$. Then, our goal is to make $\pi_{\theta}$ provide human-aligned responses for the various input prompts. To this end, we consider the popular framework of preference learning, which optimizes $\pi_{\theta}$ to learn the human preferences between two different responses [5, 18, 29]. Specifically, we assume that the preference dataset $\mathcal{D}=\{(x,y_{l},y_{w})\}$ is available which consists of the triplets of input prompt $x$, preferred response $y_{w}$, and dispreferred response $y_{l}$. Here, the preference labels were annotated by a ground truth annotator, that is usually a human expert.
Reward modeling and RL fine-tuning. Since a pairwise preference between $y_{w}$ and $y_{l}$ is hard to model directly, one of the common practices is introducing reward function $r(x,y)$ and modeling the preference based on this using the Bradley-Terry model [4]:
@@
Direct preference modeling and optimization. Rafailov et al. [32] propose an alternative approach to align LLM $\pi_{\theta}$ with preference dataset $\mathcal{D}$, which is called Direct Preference Optimization (DPO). DPO integrates a two-step alignment procedure with reward modeling and RL fine-tuning into a single unified fine-tuning procedure. Specifically, the optimal reward function is derived from RLHF objective (Eq. 3), with the target LLM $\pi_{\theta}$ and reference model $\pi_{\text{ref}}$ [12, 30, 31].
@@
4 Selfee: Boosting Alignment of LLMs with Self-generated Preference
Overview. In this section, we present Selfee: Self-generated Preference data to align LLMs while mitigating the huge cost for preference dataset construction. Our main idea is to fully exploit the human prior knowledge within the small (seed) data, and progressively update LLM to improve the alignment. To be specific, Selfee iterates two steps: (1) data expansion with self-generated preference (Section 4.1) and (2) fine-tuning LLM with self-refined preference learning (Section 4.2). We present full procedure of Selfee in Algorithm 1 (see Figure 1 for the overview).
Initial stage. We assume that a small (seed) preference dataset $D_{0}$ and an initial LLM $\pi_{\text{init}}$ are given. Here, following the common practice [29, 32, 51], we use $\pi_{\text{init}}$ which has been supervised fine-tuned (SFT) LLM on the instruction dataset [6, 43], but not aligned with human preference yet. Then, we first obtain weakly aligned LLM $\pi_{0}$ by fine-tuning $\pi_{\text{init}}$ on $D_{0}$ using DPO [32] (Eq. 6). We adopt DPO among various preference learning methods due to its simplicity and effectiveness.
Algorithm 1 Selfee algorithm
@@ 4 Selfee: Boosting Alignment of LLMs with Self-generated Preference
4.1 Data expansion with self-generated preference data
[…]
@@ 4.2 Self-refinement of generated preference data for effective learning
Self-refinement of preference label: Our key intuition is that one can view the derived preference (Eq. 5) can be viewed as the confidence of the currently training LLM $\pi_{\theta}$ for the labels assigned by $\pi_{i-1}$. Then, $\pi_{\theta}$ would exhibit lower confidence if the given pair of responses is uncertain to answer, indicating a higher probability of labeling noise. Notably, we also remark that confidence is one of the most popular metrics in the noisy label learning literature [13, 33, 36]. Under this intuition, we first identify the $K$least confident samples:
@@
where $\tau$ is the confidence of $K$ percentile sample of $\mathcal{D}_{i}$. Then, with this (potentially) noise identification label $z_{\theta}$, we refine the assigned preference label using label smoothing [25], to train $\pi_{\theta}$ less confidently when the risk of label noise is high (i.e., $z_{\theta}=1$):
@@
De-coupled noise preference detection: While learning with the refined preference label reduces the risk of learning $\pi_{\theta}$ the noisy preference, its effectiveness could be limited as the model $\pi_{\theta}$ for noise detection originated from the label generation model $\pi_{i-1}$. Therefore, to further improve the effectiveness of our denoising framework with preference label refinement, we introduce the de-coupled noise detection [13, 19] technique for LLM alignment. Specifically, we identify the preference noise by mimicking the preference prediction of a more strongly aligned LLM $\pi_{\widetilde{\theta}}$:  [^3] With $\lambda$ in Eq. 12, $\pi_{\widetilde{\theta}}$ is equivalent to model trained with (1 + $\lambda$) times smaller KL term than $\pi_{\theta}$ via Eq. 3.
@@
[…]
@@
where $\lambda>0$ is a hyper-parameter and $y_{1:t-1}$ indicates the output sequence before $t$-th output. We remark that this de-coupled noise identification by approximating $p_{\widetilde{\theta}}(y_{w}\succ y_{l}|x)$ does not require additional computations compared to DPO, since the required measurements $h_{\theta}$ and $h_{\text{ref}}$ are obtained during the calculation of the original DPO objective (Eq. 6).

### v2

# Spread Preference Annotation: Direct Preference Judgment for Efficient LLM Alignment

@@ Abstract
Aligning large language models (LLMs) with human preferences becomes a key component to obtaining state-of-the-art performance, but it yields a huge cost to construct a large human-annotated preference dataset. To tackle this problem, we propose a new framework, Spread Preference Annotation with direct preference judgment (SPA), that boosts the alignment of LLMs using only a very small amount of human-annotated preference data. Our key idea is leveraging the human prior knowledge within the small (seed) data and progressively improving the alignment of LLM, by iteratively generating the responses and learning from them with the self-annotated preference data. To be specific, we propose to derive the preference label from the logits of LLM to explicitly extract the model’s inherent preference. Compared to the previous approaches using external reward models or implicit in-context learning, we observe that the proposed approach is significantly more effective. In addition, we introduce a noise-aware preference learning algorithm to mitigate the risk of low quality within generated preference data. Our experimental results demonstrate that the proposed framework significantly boosts the alignment of LLMs. For example, we achieve superior alignment performance on AlpacaEval 2.0 with only 3.3% of the ground-truth preference labels in the Ultrafeedback data compared to the cases using the entire data or state-of-the-art baselines. [^1] https://github.com/kingdy2002/SPA
@@ 1 Introduction
Recently, large language models (LLMs) have made huge progress in various NLP tasks, leading to real-world applications that are used by millions of users, such as coding assistants and chatbots (Anthropic, 2024; OpenAI, 2022; Team et al., 2023). Aligning LLMs with human feedback, particularly through learning from human preferences, is widely considered a crucial technique for their success (Christiano et al., 2017; Lee et al., 2021; Ziegler et al., 2019). To enhance this alignment, various preference learning algorithms have been extensively explored (Ouyang et al., 2022; Rafailov et al., 2023). Despite these advancements, one of the remaining challenges is the reliance on large-scale human-annotated preference data. As the quality and quantity of the preference data are critical for the successful alignment of LLMs (Bai et al., 2022a; Cui et al., 2023), the huge cost to acquire such data inevitably presents significant obstacles.
@@
To mitigate this challenge, engaging LLMs in constructing preference data and improving their alignment using these data has recently gained attention. For example, a representative way on this line is generating multiple responses for the input prompts, and then approximating human preference between them through LLM’s predictions, often referred to as LLM-as-judge (Bai et al., 2022b; Yuan et al., 2024). However, these approaches are only effective when the given LLM is sufficiently large and well-aligned to mimic human preference via in-context learning. On the other hand, using an external reward model is considerable to substitute human preference annotation efficiently (Jiang et al., 2023b; Snorkel, 2024), but it is built on the availability of large human preference data and could also be ineffective if there is a distribution mismatch. Lastly, these approaches have a risk of potential labeling noise from LLMs, but this aspect has not been explored yet. Therefore, in this work, we aim to develop a method to effectively improve the alignment of LLM by overcoming these limitations but only relying on small human annotation.
@@
Figure 1: Illustration of the proposed SPA framework. SPA progressively improves the alignment of LLMs by iterating (1) the generation of new preference data and (2) the preference learning on the constructed data with self-refinement. Technical details are presented in Section 4.
@@
Contribution. We introduce a simple yet effective framework, coined SPA, to improve the alignment of LLMs with only a small amount of human-labeled preference data, by Spreading Preference Annotation via direct preference judgment. Our key idea is to progressively expand the knowledge of human preference within the small (seed) data, by iteratively generating the responses and learning from them through the self-annotated preference labels. Specifically, our technical contributions are three-fold as described in what follows. First, we judge the preference labels directly using the logits of LLM to explicitly extract the model’s inherent preference. This approach is more effective than previous methods that rely on external reward models or implicit in-context learning. Second, we introduce a confidence-based refinement of preference labels to reduce the risk of noise in preference learning with generated data. Third, to further enhance the effectiveness of this refinement, we propose using a linearly extrapolated prediction between current and reference models; it approximates predictions of a more strongly aligned model, leading to better noise identification.
@@
Figure 2: Summary of main result. Evaluation results on AlpacaEval 2.0 (Li et al., 2023a). Our framework significantly improves the alignment of LLMs, without additional human preference data. See detailed results in Section 5.
@@
We demonstrate the effectiveness of the proposed SPA by aligning recent LLMs with small human-annotated preference data and evaluating their alignment on the commonly used benchmarks. For example, using only 3.3% of ground-truth preference in Ultrafeedback data (Cui et al., 2023) with the mistral-7b-0.1v SFT model (Jiang et al., 2023a), our framework achieves over 16.4% increase in AlpacaEval2.0 (Li et al., 2023a) win rate compared to the initial SFT model (see Figure 2). Additionally, the AlpacaEval 2.0 length-controlled win rate is improved from 7.58% to 15.39%, and MT-bench score (Zheng et al., 2023) increased from 6.38 to 6.94. Compared to preference judgment methods like LLM-as-judge (Zheng et al., 2023), and even strong reward models such as PairRM (Jiang et al., 2023b), which have recently shown state-of-art performance in AlpacaEval2.0 benchmark, our approach consistently outperforms them across all metrics. More interestingly, the proposed SPA successfully improves the alignment of various LLMs, even without the initial human preference data. These results demonstrate that our framework is highly competitive and practical for real-world applications.
@@ 2 Related Work
Alignment of LLMs with human preference. Learning from human preferences now serves as a core component for the state-of-the-art LLMs (Anthropic, 2024; OpenAI, 2023; Team et al., 2023; Touvron et al., 2023) for aligning their responses with users’ intent and values (Ouyang et al., 2022; Ziegler et al., 2019). Arguably, one of the most popular frameworks is reinforcement learning with human preference (RLHF) (Christiano et al., 2017; Lee et al., 2021), which first trains the reward model, and then fine-tunes LLM to maximize that reward with KL divergence regularization to prevent the reward over-optimization of LLM. On the other hand, various preference learning algorithms have recently been proposed to fine-tune LLMs with human preference more efficiently (Ethayarajh et al., 2024; Hong et al., 2024; Liu et al., 2023; Rafailov et al., 2023; Xu et al., 2023; Zhao et al., 2023; Meng et al., 2024). For example, Rafailov et al. (2023) proposes Direct Preference Optimization (DPO) which allows one to fine-tune LLMs without a separate reward modeling stage, by deriving the training objective mathematically equivalent to RLHF. Ethayarajh et al. (2024) further removes the reliance on pair-wise preference labels by formulating the objective based on a human utility model. However, these methods assume that large human-annotated preference data is available, which requires a huge data acquisition cost.
Engagement of LLMs for constructing preference data. For an efficient and scalable alignment procedure, engaging LLMs for preference dataset construction has recently received attention. One common approach involves generating multiple responses to input prompts from LLM, and using an LLM’s predictions to approximate human preferences between them, a technique often referred to as LLM-as-judge (Bai et al., 2022a; Yuan et al., 2024). However, this method is effective only when the LLM is sufficiently large and well-aligned to mimic human preferences through in-context learning. Alternatively, employing an external reward model can efficiently replace human preference judgment (Jiang et al., 2023b; Snorkel, 2024), but this approach relies on the availability of extensive human preference data to pre-train reward model and may be ineffective if there is a distribution mismatch. Some recent works (Rosset et al., 2024; Snorkel, 2024; Wu et al., 2024; Xiong et al., 2024) have proposed the alignment procedure with iterative data expansion and preference learning. However, they use the external reward model or stronger LLM for the preference judgment. In contrast, we only utilize the intrinsic knowledge of training LLM for new data expansion and preference learning. Chen et al. (2025) concurrently propose a similar idea focusing on length regularization, while our work focuses on improving the quality of generated data through the proposed self-refinement procedure.
@@ 3 Preliminaries
Let us denote LLM as $\pi_{\theta}$, which generates an output sequence (e.g., response) $y$ for a given input sequence (e.g., prompt) $x$, i.e., $y\sim\pi_{\theta}(\cdot|x)$. Then, our goal is to make $\pi_{\theta}$ provide human-aligned responses to various input prompts. To this end, we consider the popular framework of preference learning, which optimizes $\pi_{\theta}$ to learn human preferences between two different responses (Christiano et al., 2017; Lee et al., 2021; Ouyang et al., 2022). Specifically, we assume that the preference dataset $\mathcal{D}=\{(x,y_{l},y_{w})\}$ is available which consists of the triplets of input prompt $x$, preferred response $y_{w}$, and dispreferred response $y_{l}$. Here, the preference labels were annotated by a ground truth annotator, that is usually a human expert.
Reward modeling and RL fine-tuning. Since a pairwise preference between $y_{w}$ and $y_{l}$ is hard to model directly, one of the common practices is introducing reward function $r(x,y)$ and modeling the preference based on this using the Bradley-Terry model (Bradley & Terry, 1952):
@@
Direct preference modeling and optimization. Rafailov et al. (2023) propose an alternative approach to align LLM $\pi_{\theta}$ with the preference dataset $\mathcal{D}$, which is called Direct Preference Optimization (DPO). DPO integrates a two-step alignment procedure with reward modeling and RL fine-tuning into a single unified fine-tuning procedure. Specifically, the optimal reward function is derived from the RLHF objective (Eq. 3), with the target LLM $\pi_{\theta}$ and the reference model $\pi_{\text{ref}}$ (Go et al., 2023; Peng et al., 2019; Peters & Schaal, 2007).
@@
4 SPA: Spread Preference Annotation to boost Alignment of LLMs
Overview. In this section, we present SPA: Spread Preference Annoation via direct preference judgment to align LLMs while mitigating the huge cost for preference dataset construction. Our main idea is to fully exploit the human prior knowledge within the small (seed) data, and progressively update LLM to improve the alignment. To be specific, SPA iterates two steps: (1) data expansion with self-generated preference (Section 4.1) and (2) fine-tuning LLM with self-refined preference learning (Section 4.2). See Figure 1 for the overview.
Initial stage. We assume that a small (seed) preference dataset $D_{0}$ and an initial LLM $\pi_{\text{init}}$ are given. Here, following the common practice (Ouyang et al., 2022; Rafailov et al., 2023; Ziegler et al., 2019), we use $\pi_{\text{init}}$ which has been supervised fine-tuned (SFT) LLM on the instruction dataset (Chung et al., 2024; Wei et al., 2022a), but not aligned with human preference yet. Then, we first obtain weakly aligned LLM $\pi_{0}$ by fine-tuning $\pi_{\text{init}}$ on $D_{0}$ using DPO (Rafailov et al., 2023) (Eq. 6). We adopt DPO among various preference learning methods due to its simplicity and effectiveness.
4.1 Direct preference judgment to align LLMs with self-generated data
[…]
Algorithm 1 SPA algorithm
@@ 4.2 Self-refinement of generated preference data for effective learning
Self-refinement of preference label: Our key intuition is that one can view the derived preference (Eq. 5) can be viewed as the confidence of the currently training LLM $\pi_{\theta}$ for the labels assigned by $\pi_{i-1}$. Then, $\pi_{\theta}$ would exhibit lower confidence if the given pair of responses is uncertain to answer, indicating a higher probability of labeling noise. Notably, we also remark that confidence is one of the most popular metrics in the noisy label learning literature (Han et al., 2018; Reed et al., 2014; Sohn et al., 2020). Under this intuition, we first identify the $K$least confident samples:
@@
where $\tau$ is the confidence of $K$ percentile sample of $\mathcal{D}_{i}$. Then, with this (potentially) noise identification label $z_{\theta}$, we refine the assigned preference label using label smoothing (Müller et al., 2019), to train $\pi_{\theta}$ less confidently when the risk of label noise is high (i.e., $z_{\theta}=1$):
@@
De-coupled noise preference detection: While learning with the refined preference label reduces the risk of learning $\pi_{\theta}$ the noisy preference, its effectiveness could be limited as the model $\pi_{\theta}$ for noise detection originated from the label generation model $\pi_{i-1}$. Therefore, to further improve the effectiveness of our preference label refinement framework, we introduce the de-coupled noise detection (Han et al., 2018; Li et al., 2020) technique for LLM alignment. Specifically, we identify the preference noise by mimicking the preference prediction of a more strongly aligned LLM $\pi_{\widetilde{\theta}}$:  [^4] With $\lambda$ in Eq. 12, $\pi_{\widetilde{\theta}}$ is equivalent to model trained with (1 + $\lambda$) times smaller KL term than $\pi_{\theta}$ via Eq. 3.
@@
With this de-coupled identification, $\pi_{\theta}$ is trained with refined preference labels via Eq. 10 , i.e., $z_{\widetilde{\theta}}$ is used to substitute $z_{\theta}$ in Eq. 10. Here, we obtain the prediction of $\pi_{\widetilde{\theta}}$ by approximating its logit $h_{\widetilde{\theta}}$ through the linear combination of the logits of $\pi_{\theta}$ and $\pi_{\text{ref}}$.  [^5] When $\pi_{\theta}(\cdot|x):=\text{Softmax}\big{(}h_{\theta}(x)\big{)}$, we refer $h_{\theta}(x)$ as the logit of LLM $\pi_{\theta}$ for the given input $x$. It is motivated by the recent work (Liu et al., 2024) that shows the aligned models via RLHF with varying $\beta$ are geometric mixtures of a reference model and a single aligned model:
@@
where $\lambda>0$ is a hyper-parameter and $y_{1:t-1}$ indicates the output sequence before $t$-th output.
We remark that this de-coupled noise identification by approximating $p_{\widetilde{\theta}}(y_{w}\succ y_{l}|x)$ does not require additional computations compared to DPO, since the required measurements $h_{\theta}$ and $h_{\text{ref}}$ are obtained during the calculation of the original DPO objective (Eq. 6). Therefore, SPA only requires a few lines of additional code to the original DPO codebase. We present full procedure of SPA in Algorithm 1.